OpenStudy (anonymous):

Solve: −3x − 1 = 8 A. −3 B. 3 C. −2.3 D. 6

OpenStudy (barrycarter):

Hint: start by adding 1 to both sides

OpenStudy (anonymous):

add one to both sides to get -3x by itself then what ever you get when you add 1 to 8, divide that by -3 to get x by itself.

OpenStudy (anonymous):

A. you would add one to each side -3x-1=8 would then go to -3x=9 then divide each side by -3 and you get x= -3

OpenStudy (anonymous):

-3x = 1 + 8 -3x = 9 x = -3